OpenStudy (anonymous):

Given log10 2=.3010 and log10 3=.4771, Find each logarithm without using a calculator. log10 square root 30

OpenStudy (anonymous):

\[\log_{10} 2=.3010 and \log_{10} 3=.4771\]

OpenStudy (anonymous):

\[\log_{10} \sqrt{30}\]

OpenStudy (anonymous):

\(30=3\times 10\) so \[\log(30)=\log(3\times 10)=\log(3)+\log(10)=\log(3)+1\]

OpenStudy (anonymous):

Well it's \[\sqrt{30}\] so then?

OpenStudy (anonymous):

\[\log(\sqrt{30}=\log(30^{\frac{1}{2}})=\frac{1}{2}\log(30)\]

OpenStudy (anonymous):

OH!!!!!! Thank you!!!!

OpenStudy (anonymous):

yw